Excerpts from Huckabee Company's December 31, 2018 and 2017, financial statements are presented below: 2018 2017 Accounts receivable $ 94,000 $ 86,000 Merchandise inventory 72,000 86,000 Net sales 414,000 386,000 Cost of goods sold 254,000 234,000 Huckabee's 2018 receivables turnover (rounded to 2 decimal places) is:
Answer #1

Average AR=(Beginning AR+Ending AR)/2

=(94000+86000)/2=$90,000

AR turnover=Credit sales/Average AR

=(414000/90000)

=4.6 times